OpenStudy (anonymous):

(Integral from 1 to infinity) [ln(x)]/[x^2] Convergent or Divergent

OpenStudy (anonymous):

\[\int\limits_{0}^{\infty} \frac{ \ln(x) }{ x ^{2} }dx\]

OpenStudy (anonymous):

i let u = lnx du = 1/x dx and dv = \[x^{-2}\] and then v = -x right??????

OpenStudy (anonymous):

First, you can compute the integral by using integration by parts. If you use int udv = uv - int vdu and choose ln x as u, you should be fine.

OpenStudy (anonymous):

im stuck on v, after that i can compute the improper integral

OpenStudy (anonymous):

dv = x^(-2) dx (don't forget the dx). Also, you're having a little trouble with v. First, get the right integral for v, you're not right on that yet.

OpenStudy (anonymous):

With getting integrals on the form x^n, you increase n by 1. That hint should get you going on the right track.

OpenStudy (anonymous):

so -x^-1???

OpenStudy (anonymous):

Yes, you're getting it now. Good job. Continue on with the parts aspect and ask a question if you get stuck.

OpenStudy (anonymous):

ok thankyou

OpenStudy (anonymous):

Once you get the indefinite integral, you can use L'Hopital's rule for limits. That will give you your answer.

OpenStudy (anonymous):

\[-x ^{-1} - \int\limits_{1}^{\infty} -x ^{-1}(\frac{ 1 }{ x })dx\]

OpenStudy (anonymous):

correct so far?

OpenStudy (anonymous):

i mean -x^-1(lnx) then integral

OpenStudy (anonymous):

Yes. I was writing a response pointing out the error that you caught yourself. You're still on the right track.

OpenStudy (anonymous):

do you have to do another integration by parts for the -x^-1(1/x) inside the integral?

OpenStudy (anonymous):

do you mind just showing me the steps? it takes a while for me to use the equation thing on here and type all of it out

OpenStudy (anonymous):

You don't have to do a second integration by parts, there you are just simplifying exponential powers of x. That last integral will be very easy. Just do exponent arithmetic and you will be doing integration on x to a certain power. Tell me if that is a big enough hint or if you still need the steps on that.

OpenStudy (anonymous):

it just simplifies to -1/x^2 right

OpenStudy (anonymous):

Yes, and then just do the integral. Remember, the second half of the formula is - int v du. So, with two negative signs, you have, as the second term, adding the int x^(-2) dx. So, watch the +- sign

OpenStudy (zarkon):

are you looking to just check for convergence/divergence as your original post indicates?

OpenStudy (anonymous):

yes but i have to show why its convergent/divergent

OpenStudy (zarkon):

\[\ln(x)<\sqrt{x}\]

OpenStudy (zarkon):

use a comparison test

OpenStudy (anonymous):

so it is converent then

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

cool, i still have to evaluate the integral

OpenStudy (zarkon):

a very simple integral

OpenStudy (anonymous):

No, you can't just do the comparison test. You have to get the indefinite integral and apply the limits on the answer to the integral.

OpenStudy (anonymous):

thats given me more problems than any other problem

OpenStudy (anonymous):

im just rusty, were doing laplace transform and havent done improper integrals in a while

OpenStudy (zarkon):

\[0\le\int\limits_{1}^{\infty} \frac{ \ln(x) }{ x ^{2} }dx\le \int\limits_{1}^{\infty} \frac{ \sqrt{x} }{ x ^{2} }dx\] \[=\int\limits_{1}^{\infty} x^{-3/2}dx=\cdots\]

OpenStudy (anonymous):

I = -x^(-1)(lnx + 1). Using L'Hopital and getting the limit of the derivative of the numerator over the denominator will show that lim as x-> inf of (ln x)/x goes to zero.

OpenStudy (anonymous):

Eventually, you will get as the integral, -x^(-1)(lnx + 1). I know you don't need the exact value for the integral, but the evaluation will give you an actual finite value which will then answer your question on convergence/divergence.

OpenStudy (anonymous):

im sorry i just really am confused now... before zarkon came in i got to \[-\infty ^{-1}(\ln \infty) - (-0^{-1}(\ln 0) + -\frac{ 1 }{ 3 }x ^{-3}\]

OpenStudy (anonymous):

and im not confident in that at all

OpenStudy (anonymous):

i have to evaluate the -1/3x^-3 from infinity - 0 too

OpenStudy (anonymous):

Don't worry, I'll get you through the rest of it.

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

First, did you get as the final integral -x^(-1)(lnx + 1)?

OpenStudy (anonymous):

no i just got -x^-1(lnx)

OpenStudy (anonymous):

i never had a +1 in mine

OpenStudy (zarkon):

are you going from 1 to infinity or 0 to infinity?

OpenStudy (anonymous):

1 to infinity.. i accidentally put 0 in the opening equation

OpenStudy (anonymous):

The 1 is from doing - int v du. I'll show you the steps. You've worked way hard on this and you have done well, so I'm not really giving away the store here. You did a tremendous amount of work.

OpenStudy (zarkon):

ok...carry on ;)

OpenStudy (anonymous):

I = -x^(-1)lnx + int x^(-2) dx. the integration by parts is uv - int v du, but you have a negative in v, so there's a double negative going on there that I just combined.

OpenStudy (anonymous):

i never set it to equal 1 like that

OpenStudy (anonymous):

i have that on my paper

OpenStudy (anonymous):

That -x^(-1)lnx (first term) is uv. So, I = -x^(-1)lnx - x^(-1) or -x^(-1)(ln x + 1). Limit of that as x -> inf is the limit of the derivative of the numerator over the derivative of the denominator (L'Hopital). So, limit of (-1/x)/1 as x -> inf, so that would be 0. So, all you have to do is evaluate at 1 (subtracting, again a double minus). So, x^(-1)(ln x + 1) (notice no negative now as we subtracted a negative) at 1 is 1^(-1)[(ln 1) + 1] = 1. Convergent.

OpenStudy (anonymous):

Sweet thank you for all your help, I really appreciate that

OpenStudy (anonymous):

I hope I was helping. It's hard to show so many steps and not lose someone. I really do hope this has lifted a burden for you and that you get what I was trying to say. If not, I blame myself as teacher. You obviously are a good and hard-working student. Good luck to you.

OpenStudy (anonymous):

I completely get what you were trying to say. I got lost there for a second but once you typed out the final steps it made sense. thanks again

OpenStudy (anonymous):

You have a wonderful day. and you be a good engineer or whatever!